OpenStudy (anonymous):

The ratio of the areas of two similar sundeck porches is 225:36. If the larger sundeck is 15 feet along one side, how long is the smaller sundeck along its corresponding side? I don't want anyone to do it for me but to give me the steps so I could try to do it myself.

OpenStudy (anonymous):

its easy u have formula folow that .. that will make ur concept........ a:b::x:y solve it as a*y=b*x

OpenStudy (anonymous):

so it is 225:36 and 15:y 225(y)=36(15) 225y=540 y=540/225 y=2.4

OpenStudy (anonymous):

u got it .......

OpenStudy (anonymous):

ok. thanks

OpenStudy (anonymous):

mention not